Shadee Corp. expects to sell 510 sun visors in May and 420 in June. Each visor sells for $21. Shadee’s beginning and ending finished goods inventories for May are 80 and 45 units, respectively. Ending finished goods inventory for June will be 70 units.

Suppose that each visor takes 0.30 direct labor hours to produce and Shadee pays its workers $10 per hour.

Determine Shadee's budgeted direct labor cost for May and June.
Answer #1

(1)

For May

budgeted production of visors = budgeted sales + ending inventory - beginning inventory

= 510 + 45 - 80 = 475

budgeted direct labor cost = budgeted production x direct labor hour per unit produced x rate per hour

= 475 x 0.30 x $10

= $1425

(2)

For June,

budgeted production of visors = budgeted sales + ending inventory - beginning inventory

= 420 + 70 - 45 = 445

budgeted direct labor cost = budgeted production x direct labor hour per unit produced x rate per hour

= 445 x 0.30 x $10

= $1335